# OmniNav: A Unified Framework for Prospective Exploration and Visual-Language Navigation

### v1

@@ Abstract
Embodied navigation is a foundational challenge for intelligent robots, demanding the ability to comprehend visual environments, follow natural language instructions, and explore autonomously. However, existing models struggle to provide a unified solution across heterogeneous navigation paradigms, often yielding low success rates and limited generalization. We present OmniNav, a unified framework that handles instruct-goal, object-goal, point-goal navigation, and frontier-based exploration within a single architecture. First, we introduce a lightweight, low-latency policy that predicts continuous-space waypoints (coordinates and orientations) with high accuracy, outperforming action-chunk methods in precision and supporting real-world deployment with control frequencies up to 5 Hz. Second, at the architectural level, OmniNav proposes a fast-slow system design: a fast module performs waypoint generation from relatively short-horizon visual context and subtasks, while a slow module conducts deliberative planning using long-horizon observations and candidate frontiers to select the next subgoal and subtask. This collaboration improves path efficiency and maintains trajectory coherence in exploration and memory-intensive settings. Notably, we find that the primary bottleneck lies not in navigation policy learning per se, but in robust understanding of general instructions and objects. To enhance generalization, we incorporate large-scale general-purpose training dataset including those used for image captioning and visual into a joint multi-task regimen, which substantially boosts success rates and robustness. Extensive experiments demonstrate state-of-the-art performance across diverse navigation benchmarks, and real-world deployment further validates the approach. OmniNav offers practical insights for embodied navigation and points to a scalable path toward versatile, highly generalizable robotic intelligence.
@@ 1 INTRODUCTION
Embodied navigation (Gao et al., 2024; Gu et al., 2022) has emerged as a core problem in embodied intelligence: enabling robots to perceive, understand, and explore real-world environments without pre-built maps while following natural language instructions. To act reliably in dynamic, partially observable environments, an agent must not only ground instantaneous visual inputs but also maintain coherent spatiotemporal memory and perform active exploration. Application demands for real-time responsiveness further increase the requirements for low-latency decision-making and cross-environment generalization.
@@
Current research largely revolves around three paradigms: point-goal (Liu et al., 2025), instruct-goal (Anderson et al., 2018; Ku et al., 2020), and object-goal (Yokoyama et al., 2024b). point-goal tasks are well-specified and straightforward to evaluate but rely on explicit coordinates rarely available in practice; instruction-goal aligns with human usage but often generalizes poorly to unseen instructions or environments; object-goal is the most practical but requires robust target recognition coupled with efficient path planning, making it the most challenging. Many existing methods remain customized, relying on task-specific data, which limits cross-task transfer and the potential for mutual enhancement. Uni-Navid (Zhang et al., 2024a) proposed a VLM-based discrete action predictor unifying vision-and-language navigation, object-goal navigation, embodied question answering  (Das et al., 2018), and following (Wang et al., 2025a), but its study of LLM long-horizon planning is not sufficiently developed. MTU3D (Zhu et al., 2025) advances a “move to understand” paradigm by coupling frontier exploration with visual localization in a single objective, yet requires constructing 3D object coordinates, leading to deployment complexity. Although recent Video-LLMs (Wei et al., 2025; Qi et al., 2025; Zhang et al., 2024b) and VLAs (Sapkota et al., 2025; Zitkovich et al., 2023; Ma et al., 2024) integrate vision, language, and action prediction end to end, they still face bottlenecks in streaming video input, long-context management, and low-latency inference: discretized action modeling sacrifices precision and flexibility; constrained LLM call frequency and frequent context resets lead to deployment difficulties; besides, in practice, the dominant failure mode often stems from inadequate understanding of generic instructions and open-vocabulary objects rather than policy learning itself. These gaps call for a unified, efficient framework that balances long/short-horizon reasoning with real-time responsiveness.
@@
We present OmniNav, a unified embodied navigation framework that concurrently covers instruct-goal, object-goal, point-goal, and frontier-based exploration within a single architecture. Inspired by dual-system theory (Figure, 2024; Black et al., 2025), OmniNav coordinates a fast–slow system (Black et al., 2025): a fast system reacts to comparatively short-horizon perception and current tasks or subtasks, generating high-precision waypoints (coordinates and orientations) to support low-latency control up to 5 Hz; a slow system deliberates over long-horizon observations and frontier cues, leveraging a VLM’s chain-of-thought (Wei et al., 2022) to decompose complex goals and select the next subgoal and subtask. The two are coupled through a central memory module that uses a key–value (KV) cache to provide essential spatiotemporal context, yielding decisions that are both locally agile and globally consistent.
@@
OmniNav addresses the triad of real-time operation, fast–slow collaboration, and generalization. A lightweight flow-matching policy (Bjorck et al., 2025) avoids the precision degradation and latency accumulation inherent to action discretization; fast–slow collaboration ensures exploration efficiency and trajectory coherence in long-memory scenarios; and, more importantly, training unifies large-scale generic vision–language data (captioning, referring/grounding, etc.) with multiple navigation tasks, significantly strengthening instruction following and open-vocabulary object perception to improve success rates and robustness. Our contributions are threefold:
  • The first unified architecture to simultaneously support instruction, object, and point goals alongside frontier exploration;
@@ 2 RELATED WORKS
Vision Language Models for Navigation Leveraging their powerful generalization capabilities in understanding and planning, Visual Language Models (VLMs) (Chiang et al., 2023; Liu et al., 2023; Zhu et al., 2023) have been increasingly applied to the domain of robotic navigation, achieving notable success. Prevailing methods (Dorbala et al., 2022; Zhou et al., 2024b; Long et al., 2024b) typically employ VLMs to process multimodal instructions and directly decode low-level actions in an autoregressive manner. However, this paradigm suffers from significant drawbacks: it is prone to compounding errors in sequential prediction and is often hampered by slow inference speeds. In contrast, our approach draws inspiration from recent advances in Vision-Language-Action (VLA) models (Zitkovich et al., 2023; Kim et al., 2024; Li et al., 2024). We introduce a novel architecture that appends a flow-matching policy (Zhao et al., 2024; Chen et al., 2024; Zhou et al., 2024a) to a VLM backbone. This design enables our model to generate entire action trajectories non-autoregressively, leading to substantially improved prediction accuracy and computational efficiency, especially when navigating unseen environments.
Dual-System Design Dual-system architectures have been widely adopted across various domains to meet diverse operational demands. In the realm of Vision-Language-Action (VLA) models, several works (Bjorck et al., 2025; Bu et al., 2025; Ge et al., 2024; Song et al., 2025) have implemented dual-system designs to balance fast control execution and intelligent planning. Inspired by this paradigm and motivated by the specific requirements of embodied navigation, we propose a novel dual-system framework. Our framework consists of two complementary components. The first is a fast system, a purely visual, end-to-end policy designed for direct deployment and is highly effective in the majority of navigation scenarios. The second system is specifically engineered for challenging long-horizon tasks. It leverages a frontier-based planning strategy (Zhu et al., 2025) to explicitly address the limitations of finite context windows and prevent performance degradation when faced with extensive observation histories. This synergistic design has proven its superiority by achieving State-of-the-Art performance on multiple established benchmarks.
@@ 3 Approach
Figure 1: The fast thinking system can independently handle multi-task navigation, using a VLM backbone and a flow-matching policy to rapidly generate waypoints. Building on this, a slow thinking module is integrated to enable long-term memory and planning: it constructs long-range spatial and semantic memory using frontiers and images, and provides subgoal cues. The reasoning process is briefly summaried as: if the target already exists in memory, the fast thinking module is invoked to reach it; otherwise, the system selects the most appropriate subgoal for next exploration.
@@
Multimodal Input tokenizations To handle all four task types through a unified interface, text, coordinates, and visual history are converted into a set of discrete tokens consumable by a Large-Language Model (LLM) see Fig. 1. We use Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as the base model and extend it with a coordinate modality. During streaming inference, a key–value (KV) cache is maintained to reduce latency. Text tokens: Derived from natural-language task descriptions, object category labels, and point-goal commands, are all converted into a standardized instruction sequence. Coordinate tokens: Candidate search regions are represented as sets of 2D coordinates and heading angles, sourced from point-goal inputs or subgoal positions generated by the slow system. These coordinates are processed via an MLP to dense embeddings that serve as coordinate tokens. Image tokens: The central memory maintains a ring buffer of pose-stamped images. For the fast system, it spatiotemporally samples from the historical image sequence, maintaining a maximum number of images (e.g., 20 frames). For the slow system, it samples images from the spatiotemporal neighborhood of candidate frontiers. All images are encoded with a ViT to produce image tokens.
@@
Fast Thinking System OmniNav operates at a high frequency, designed to execute either subtasks provided by a slow system or end-to-end multi-task navigation, as shown in Fig. 1. It directly outputs a sequence of continuous-space waypoints $\mathbf{w}_{t:t+H}\in\mathbb{R}^{H\times 5}$ with $H=5$, enabling real-time closed-loop control. We formulate waypoint prediction as a conditional diffusion generation, interacted with the fused features from the vision-language model (VLM). This design preserves the VLM’s semantic understanding while enabling rich interactions between language-guided context and the waypoint, leading to robust instruction following. Compared to discrete action chunking, our approach produces smoother, higher-precision trajectories with lower inference latency, supporting stable 5 Hz execution on hardware. We employ a variant of the Denoising Transformer (DiT) (Peebles & Xie, 2023) to model waypoint sequences. The policy network consists of self-attention blocks that operate on noised tokens to capture temporal and spatial dependencies within the waypoint sequence, and cross-attention blocks that attend to the vision-language context $\mathbf{O}_{VLM}$. The output is a sequence of $H=5$ spatial-temporal waypoints $\mathbf{w}_{t}^{(i)}\in\mathbb{R}^{5}$, $i=1,\dots,H$, each encoding:
@@
Conditional flow matching policy is employed (Lipman et al., 2022). Given a ground-truth waypoint sequence $\mathbf{w}_{t:t+H}$, noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and a time parameter $\tau\in[0,1]$, the input is constructed as:
$\mathbf{w}_{t:t+H}^{\tau}=\tau\mathbf{w}_{t:t+H}+(1-\tau)\bm{\epsilon},$ (2)
and the policy $\pi$ is trained to estimate the denoising residual $\bm{\epsilon}-\mathbf{w}_{t:t+H}$ by minimizing:
$\mathbb{E}_{\tau,\bm{\epsilon}}\left[\left\|\pi(\mathbf{O}_{VLM},\mathbf{w}_{t:t+H}^{\tau})-(\bm{\epsilon}-\mathbf{w}_{t:t+H})\right\|^{2}\right].$ (3)
@@
Slow Thinking System The slow thinking system is the deliberative planning module for hierarchical active exploration. It publishes executable spatial subgoals and subtasks to the fast thinking system. Its core responsibilities are twofold: When the target appears in the current or historical field of view, it can quickly localize the target and generate subgoal and subtask that drive the fast system to progressively approach it. When the target is not observed, it selects subgoal position with strong semantic relevance to the target to explore next. This process demands both exploration and environmental understanding: the model must grasp spatial locations as well as semantic information. Frontier (Zhu et al., 2025) is employed to guide the active exploration. We maintain a 3D occupancy map, which categorizes each region as explored or unknown, and frontiers are then identified as the boundary points between explored and unknown regions. In addition, to comprehend past temporal and spatial information, we construct a memory bank (Zhu et al., 2025; Olton, 1977; Xu et al., 2024; Zhou et al., 2023). This repository archives a history of observations, storing the visual data and corresponding pose information (coordinates and orientations) after every executed action. We then design a sampling strategy that connects this historical context to future exploration by collecting all historical images captured near the agent’s current location. It then evaluates each frontier by iterating through these images, sampling the one whose original capture viewpoint is most suitably aligned with the frontier’s spatial coordinates. This image thereby becomes a visual proxy for that frontier.
@@
During frontier selection, the model engages in comprehensive spatial and content reasoning, reflecting its capability to actively explore unknown environments and propose subgoal locations related to the target object. For example, when searching for a toilet, it prioritizes exploration locations in the bathroom; when searching for a television, it seeks locations associated with the living room. Moreover, if the target object appears in memory or within the current view, it outputs the target’s existing location. We incorporate explicit chain-of-thought (CoT) (Lin et al., 2025; Wang et al., 2025b; Zhao et al., 2025) reasoning into the slow system’s prediction process to enable transparent process expression, achieving interpretability and self-correction. This also allows for richer textual outputs that strengthen the model’s grasp of logic and improve complex reasoning performance. Fig.2 illustrates the slow system’s reasoning process.

### v3

@@ Abstract
Embodied navigation is a foundational challenge for intelligent robots, demanding the ability to comprehend visual environments, follow natural language instructions, and explore autonomously. However, existing models struggle to provide a unified solution across heterogeneous navigation paradigms, often yielding low success rates and limited generalization. We present OmniNav, a unified framework that handles instruct-goal, object-goal, point-goal navigation, and frontier-based exploration within a single architecture. First, we introduce a lightweight, low-latency policy that predicts continuous-space waypoints (coordinates and orientations) with high accuracy, outperforming action-chunk methods in precision and supporting real-world deployment with control frequencies up to 5 Hz. Second, at the architectural level, OmniNav proposes a fast-slow system design: a fast module performs waypoint generation from relatively short-horizon visual context and subtasks, while a slow module conducts deliberative planning using long-horizon observations and candidate frontiers to select the next subgoal and subtask. This collaboration improves path efficiency and maintains trajectory coherence in exploration and memory-intensive settings. Notably, we find that the primary bottleneck lies not in navigation policy learning per se, but in robust understanding of general instructions and objects. To enhance generalization, we incorporate large-scale general-purpose training datasets including those used for image captioning and referring/grounding into a joint multi-task regimen, which substantially boosts success rates and robustness. Extensive experiments demonstrate state-of-the-art performance across diverse navigation benchmarks, and real-world deployment further validates the approach. OmniNav offers practical insights for embodied navigation and points to a scalable path toward versatile, highly generalizable robotic intelligence.
@@ 1 INTRODUCTION
Embodied navigation (gao2024vision; gu2022vision) has emerged as a core problem in embodied intelligence: enabling robots to perceive, understand, and explore real-world environments without pre-built maps while following natural language instructions. To act reliably in dynamic, partially observable environments, an agent must not only ground instantaneous visual inputs but also maintain coherent spatiotemporal memory and perform active exploration. Application demands for real-time responsiveness further increase the requirements for low-latency decision-making and cross-environment generalization.
@@
Current research largely revolves around three paradigms: point-goal (liu2025citywalker), instruct-goal (anderson2018vision; ku2020room), and object-goal (yokoyama2024hm3d). point-goal tasks are well-specified and straightforward to evaluate but rely on explicit coordinates rarely available in practice; instruction-goal aligns with human usage but often generalizes poorly to unseen instructions or environments; object-goal is the most practical but requires robust target recognition coupled with efficient path planning, making it the most challenging. Many existing methods remain customized, relying on task-specific data, which limits cross-task transfer and the potential for mutual enhancement. Uni-Navid (zhang2024uni) proposed a VLM-based discrete action predictor unifying vision-and-language navigation, object-goal navigation, embodied question answering (das2018embodied), and following (wang2025trackvla), but its study of LLM long-horizon planning is not sufficiently developed. MTU3D (zhu2025move) advances a “move to understand” paradigm by coupling frontier exploration with visual localization in a single objective, yet requires constructing 3D object coordinates, leading to deployment complexity. Although recent Video-LLMs (wei2025streamvln; qi2025vln; zhang2024navid) and VLAs (sapkota2025vision; zitkovich2023rt; ma2024survey) integrate vision, language, and action prediction end to end, they still face bottlenecks in streaming video input, long-context management, and low-latency inference: discretized action modeling sacrifices precision and flexibility; constrained LLM call frequency and frequent context resets lead to deployment difficulties; besides, in practice, the dominant failure mode often stems from inadequate understanding of generic instructions and open-vocabulary objects rather than policy learning itself. These gaps call for a unified, efficient framework that balances long/short-horizon reasoning with real-time responsiveness.
@@
We present OmniNav, a unified embodied navigation framework that concurrently covers instruct-goal, object-goal, point-goal, and frontier-based exploration within a single architecture. Inspired by dual-system theory (figure2024helix; black2025pi0), OmniNav coordinates a fast–slow system (black2025pi0): a fast system reacts to comparatively short-horizon perception and current tasks or subtasks, generating high-precision waypoints (coordinates and orientations) to support low-latency control up to 5 Hz; a slow system deliberates over long-horizon observations and frontier cues, leveraging a VLM’s chain-of-thought (wei2022chain) to decompose complex goals and select the next subgoal and subtask. The two are coupled through a central memory module that uses a key–value (KV) cache to provide essential spatiotemporal context, yielding decisions that are both locally agile and globally consistent.
@@
OmniNav addresses the triad of real-time operation, fast–slow collaboration, and generalization. A lightweight flow-matching policy (bjorck2025gr00t) avoids the precision degradation and latency accumulation inherent to action discretization; fast–slow collaboration ensures exploration efficiency and trajectory coherence in long-memory scenarios; more importantly, training unifies large-scale generic vision–language data (captioning, referring/grounding, etc.) with multiple navigation tasks, significantly strengthening instruction following and open-vocabulary object perception to improve success rates and robustness. Our contributions are threefold:
  • A unified architecture that, under a single training framework and policy, supports multiple goal modalities (point, object, and instruction) as well as frontier-based exploration;
@@ 2 RELATED WORKS
Vision Language Models for Navigation Leveraging their powerful generalization capabilities in understanding and planning, Visual Language Models (VLMs) (chiang2023vicuna; liu2023visual; zhu2023chatgpt) have been increasingly applied to the domain of robotic navigation, achieving notable success. Prevailing methods (dorbala2022clip; zhou2024navgpt; long2024discuss) typically employ VLMs to process multimodal instructions and directly decode low-level actions in an autoregressive manner. However, this paradigm suffers from significant drawbacks: it is prone to compounding errors in sequential prediction and is often hampered by slow inference speeds. In contrast, our approach draws inspiration from recent advances in Vision-Language-Action (VLA) models (zitkovich2023rt; kim2024openvla; li2024manipllm). We introduce a novel architecture that appends a flow-matching policy (zhao2024monoformer; chen2024diffusion; zhou2024transfusion) to a VLM backbone. This design enables our model to generate entire action trajectories non-autoregressively, leading to substantially improved prediction accuracy and computational efficiency, especially when navigating unseen environments.
Dual-System Design Dual-system architectures have been widely adopted across various domains to meet diverse operational demands. In the realm of Vision-Language-Action (VLA) models, several works (bjorck2025gr00t; bu2025agibot; ge2024seed; song2025hume) have implemented dual-system designs to balance fast control execution and intelligent planning. Inspired by this paradigm and motivated by the specific requirements of embodied navigation, we propose a novel dual-system framework. Our framework consists of two complementary components. The first is a fast system, a purely visual, end-to-end policy designed for direct deployment and is highly effective in the majority of navigation scenarios. The second system is specifically engineered for challenging long-horizon tasks. To serve as a long-term memory mechanism, we employ a planning strategy that combines frontier-based exploration (zhu2025move) with images. This approach offers a more concise implementation compared to alternative memory structures such as scene graphs (team2025robobrain) or complex semantic maps (long2024instructnav). These alternatives memory structures can also be implementations for the slow system. The idea stays the same: the slow system is responsible for global planning, while the fast system handles local execution. This synergistic design has proven its superiority by achieving State-of-the-Art performance on multiple benchmarks.
Frontier-based Navigation Recent studies on exploration and navigation adopt different strategies for selecting informative targets in unknown environments. GOAT (chang2023goat) and its benchmark GOAT-Bench (khanna2024goat) study lifelong navigation and object search using an object-instance memory and frontier exploration. Similarly, MTU3D (zhu2025move) keep an object-goal memory built from 3D point clouds and semantic segmentation, and combine this with frontier exploration. A different group of methods uses non-semantic frontier exploration (chang2023goat; sakamoto2024map; nayak2025multi), where the next target is usually just the closest frontier, sometimes adjusted by simple heuristics such as distance–heading scores. OmniNav instead uses a semantics- and reasoning-aware frontier selection: it links each frontier to its egocentric images, then uses explicit chain-of-thought reasoning over these views to decide which frontier is more informative or promising for the current task.
@@ 3 Approach
Figure 1: The fast system can independently handle multi-task navigation, using the VLM backbone and a flow-matching policy to rapidly generate waypoints. Building on this, a slow thinking module is integrated to enable long-term memory and planning: it constructs long-range spatial and semantic memory using frontiers and images, and provides subgoal cues. The collaboration between the slow and fast proceeds as follows: the slow system uses frontiers or memory to generate high-level subgoals, once a subgoal is determined, the fast system takes over and progressively produces low-level waypoint sequences, ultimately reaching the target.
@@
Multimodal Input tokenizations To handle all four task types through a unified interface, text, coordinates, and visual history are converted into a set of discrete tokens consumable by a Large-Language Model (LLM) see Fig. 1. We use Qwen2.5-VL-3B-Instruct (bai2025qwen2) as the base model and extend it with a coordinate modality. During streaming inference, a key–value (KV) cache is maintained to reduce latency. Text tokens: Derived from natural-language task descriptions, object category labels, and point-goal commands, are all converted into a standardized instruction sequence. Coordinate tokens: Candidate search regions are represented as sets of 2D coordinates and heading angles, sourced from point-goal inputs or subgoal positions generated by the slow system. These coordinates are processed via an MLP to dense embeddings that serve as coordinate tokens. Image tokens: The central memory maintains a ring buffer of pose-stamped images. For the fast system, it spatiotemporally samples from the historical image sequence, maintaining a maximum number of images (e.g., 20 frames). For the slow system, it samples images from the spatiotemporal neighborhood of candidate frontiers. All images are encoded with a ViT to produce image tokens.
@@
Fast Thinking System OmniNav operates at a high frequency, designed to execute either subtasks provided by a slow system or end-to-end multi-task navigation, as shown in Fig. 1. It parallelly outputs a sequence of 5 continuous-space waypoints, $\mathbf{w}_{t:t+H}\in\mathbb{R}^{H\times 5}$ with $H=5$. We formulate waypoint prediction as a conditional diffusion generation task. The input coordinates are first embedded by an MLP and then encoded together with the images and texts by the VLM. The VLM performs deep fusion over these features, and the resulting fused features are used as conditions for the diffusion model, guiding waypoint generation. This design preserves the VLM’s semantic understanding while enabling rich interactions between language-guided context and the waypoint, leading to robust instruction following. Compared with conventional autoregressive methods, the policy head achieves an speedup and, with a history of 20 frames, supports an inference rate of 5 Hz for real-time closed-loop control. At the same time, it produces smoother and more precise trajectories.
We employ a variant of the Denoising Transformer (DiT) (peebles2023scalable) to model waypoint sequences. The policy network consists of self-attention blocks that operate on noised tokens to capture temporal and spatial dependencies within the waypoint sequence, and cross-attention blocks that attend to the vision-language context $\mathbf{O}_{VLM}$. The output is a sequence of $H=5$ spatial-temporal waypoints $\mathbf{w}_{t}^{(i)}\in\mathbb{R}^{5}$, $i=1,\dots,H$, each encoding:
@@
Conditional flow matching policy is employed (lipman2022flow). Given a ground-truth waypoint sequence $\mathbf{w}_{t:t+H}$, noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and a time parameter $\tau\in[0,1]$, the input is constructed as:
$\mathbf{w}_{t:t+H}^{\tau}=\tau\mathbf{w}_{t:t+H}+(1-\tau)\boldsymbol{\epsilon},$ (2)
and the policy $\pi$ is trained to estimate the denoising residual $\boldsymbol{\epsilon}-\mathbf{w}_{t:t+H}$ by minimizing:
$\mathbb{E}_{\tau,\boldsymbol{\epsilon}}\left[\left\|\pi(\mathbf{O}_{VLM},\mathbf{w}_{t:t+H}^{\tau})-(\boldsymbol{\epsilon}-\mathbf{w}_{t:t+H})\right\|^{2}\right].$ (3)
@@
Slow Thinking System The slow thinking system is the deliberative planning module for hierarchical active exploration. Its core responsibilities are twofold: when the target appears in the current or historical field of view, it can quickly localize the target and generate the coordinates of the subgoal that drive the fast system to progressively approach it, when the target is not observed, it selects subgoal position with strong semantic relevance to the target to explore next. This process demands both exploration and environmental understanding. The former requires the model to navigate and discover the environment, while the latter involves a core task: predicting the spatial coordinates of the target based on input camera poses and images, this grounds the model’s semantic understanding in a concrete, geometric output.
In the fast–slow system collaboration, the fast system is not just a low-level controller following preset coordinates or smoothing a pre-planned path. It must constantly use raw visual input to move toward its subgoal. For example, if it gets a coordinate from the slow system but a wall blocks the straight-line path, the fast system must use visual cues to find a path that avoids the obstacle. If it gets a target coordinate from object memory, it can adjust its final pose based on what it currently sees, such as stopping precisely at the left, right, or center of the object. By continuously updating and refining its waypoints while moving, the fast system can reach targets more accurately.
@@
Frontier (zhu2025move) is employed to guide the active exploration. We maintain a 3D occupancy map, which categorizes each region as explored or unknown, and frontiers are then identified as the boundary points between explored and unknown regions. In addition, to comprehend past temporal and spatial information, we construct a memory bank (zhu2025move; olton1977spatial; xu2024embodiedsam; zhou2023dual). This repository archives a history of observations, storing the visual data and corresponding pose information (coordinates and orientations) after every executed action. We then design a sampling strategy that connects this historical context to future exploration by collecting all historical images captured near the agent’s current location. It then evaluates each frontier by iterating through these images, sampling the one whose original capture viewpoint is most suitably aligned with the frontier’s spatial coordinates. This image thereby becomes a visual proxy for that frontier.
@@
During frontier selection, the model engages in comprehensive spatial and content reasoning, reflecting its capability to actively explore unknown environments and propose subgoal locations related to the target object. For example, when searching for a toilet, it prioritizes exploration locations in the bathroom; when searching for a television, it seeks locations associated with the living room. Moreover, if the target object appears in memory or within the current view, it outputs the target’s existing location. We incorporate explicit chain-of-thought (CoT) (lin2025navcot; wang2025think; zhao2025cot) reasoning into the slow system’s prediction process to enable transparent process expression, achieving interpretability and self-correction. This also allows for richer textual outputs that strengthen the model’s grasp of logic and improve complex reasoning performance. Fig.2 illustrates the slow system’s reasoning process.